OpenStudy (anonymous):

Solve the DE. e^(-t/50)-(Q/50) = Q' Q(0) = 0 I got the following answer but it is not correct: [-100e^(-t/50)+100]/[e^(-t/50)]

OpenStudy (anonymous):

I did the following: Wrote it in standard linear form: Q'(t)+(1/50)Q(t)=e^(-t/50) So my integrating factor was e^(integral(1/50)dt))=e^(1/50)t Did you do the same?

OpenStudy (anonymous):

I got (t+c)/(e^(t/50)) = Q However, it gets flagged as incorrect.

OpenStudy (anonymous):

Yeah that's correct up to that point. You can bust up that fraction and use Q(0) = 0 to solve for c